Weight in Pounds 3-4 5-6 7-8 9-10 11-12 13-14 Number of Newborns 4 Based on the frequency distribution above, find the relative frequency for the class with lower class limit 7 Relative Frequency Give your answer as a percent, rounded to one decimal place
Answer #1

RELATIVE FREQUENCY FOR THE CLASS WITH LOWER CLASS LIMIT 7 IS,

= FREQUENCY / TOTAL FREQUENCY = 4/27 = 0.148148

RELATIVE FREQUENCY IN PERCENTAGE = 0.148148 * 100 = 14.8 %